# BEV-SUSHI: Multi-Target Multi-Camera 3D Detection and Tracking in Bird’s-Eye View

### v2

@@ Abstract
Object perception from multi-view cameras is crucial for intelligent systems, particularly in indoor environments, e.g., warehouses, retail stores, and hospitals. Most traditional multi-target multi-camera (MTMC) detection and tracking methods rely on 2D object detection, single-view multi-object tracking (MOT), and cross-view re-identification (ReID) techniques, without properly handling important 3D information by multi-view image aggregation. In this paper, we propose a 3D object detection and tracking framework, named BEV-SUSHI, which first aggregates multi-view images with necessary camera calibration parameters to obtain 3D object detections in bird’s-eye view (BEV). Then, we introduce hierarchical graph neural networks (GNNs) to track these 3D detections in BEV for MTMC tracking results. Unlike existing methods, BEV-SUSHI has impressive generalizability across different scenes and diverse camera settings, with exceptional capability for long-term association handling. As a result, our proposed BEV-SUSHI establishes the new state-of-the-art on the AICity’24 dataset with $81.22$ HOTA, and $95.6$ IDF1 on the WildTrack dataset.
@@ 1 Introduction
Detecting and tracking objects across multiple cameras is a crucial problem for indoor environment understanding, particularly in retail or warehouse settings (an example is shown in Fig. 1), for various applications, including inventory management, security surveillance, or customer behavior analysis. In these use cases, a typical multi-camera system involves numerous cameras with diverse viewing angles and fields of view (FoV). While some cameras usually have overlapping FoVs, others may not share any common visual space. This configuration allows for comprehensive coverage of the scene while balancing between redundancy and efficiency. However, this task presents significant challenges such as occlusions, varying lighting conditions, and the need to maintain consistent object identification across different camera views. Moreover, the integration of 3D spatial information from multiple 2D views requires sophisticated algorithms to handle camera calibration errors and perspective distortions. Addressing these issues is crucial for developing robust multi-camera detection and tracking systems that can reliably operate in complex indoor environments.
Figure 1: Example of an indoor environment (warehouse) with 8 cameras generated by NVIDIA Omniverse Platform [22]. The world coordinates $\mathcal{W}$ and camera coordinates $\mathcal{C}^{i}$ are drawn in the top-down visualization on the bottom right. Objects with 3D bounding boxes are annotated in both BEV and camera views.
@@
Existing MTMC tracking methods can be divided into three categories: (i) late multi-view aggregation, (ii) late multi-view aggregation with geometric projection, and (iii) early multi-view aggregation, as shown in Fig. 2. Late multi-view aggregation pipelines detect objects in each camera view as 2D bounding boxes, with some methods tracking these 2D detections separately before associating them across cameras using re-identification (ReID) embeddings and spatial constraints. Late multi-view aggregation pipelines with geometric projection pipelines [34, 35, 8] further perform additional projections using camera calibration matrices to achieve spatial association for each camera view into a global coordinate system. Recently,  [29] demonstrated the possibility of early aggregation before any perception steps in each camera view. This type of method first fuses multi-view image information into a unified 3D space and directly conducts both detection and tracking in this 3D space, which can significantly improve detection quality and avoid association errors due to unreliable spatial alignment among different camera views. However, its detection network is designed for a fixed multi-camera scene, so it is not flexible for different environments, different numbers of cameras, or different camera placements. Besides, while it outputs robust 3D detections, the method then uses an online Kalman filter based tracking which is prone to drifting and struggles with occlusions.
In this work, we propose BEV-SUSHI, a generalizable method capable of robust 3D detection and long-term tracking in various complex multi-camera environments.
Our approach leverages a unified BEV representation and spatio-temporal transformers to directly infer 3D object detections, eliminating the need for 2D detections and their subsequent projection to 3D space. This method bypasses the alignment step by projecting 2D information from multiple cameras into the unified BEV coordinates, resulting in a comprehensive and accurate scene representation. We also introduce a 2D-3D detection association process for ReID embedding extraction and detection verification, further enhancing the robustness and accuracy of our system.
@@
For the temporal association of detections, we introduce the first 3D tracking-by-detection approach utilizing hierarchical graph neural networks (GNNs) [5]. Its lightweight hierarchical structure allows us to learn and model track associations for up to 4k frames. Even longer video clips are processed in a near-online sliding window fashion. Our model consists of a hierarchy of GNNs to associate 3D detections via geometrical and multi-view ReID embedding features. To improve long-term tracking, we introduce a novel global tracking layer that replaces the handcrafted tracking heuristics in [5] with an additional layer that matches incoming with past tracks. The global layer requires no additional training and significantly boosts occlusion handling and overall tracking performance.
@@
Finally, we evaluate BEV-SUSHI on a large-scale synthetic indoor MTMC dataset called AICity’24 [31] and a real-world outdoor MTMC dataset named WildTrack [6]. BEV-SUSHI achieves the SOTA on both datasets, with $81.22$ HOTA on AICity’24, and $95.6$ IDF1 on WildTrack.
@@
  • A multi-view image aggregation network to achieve 3D detection in large-scale multi-camera environments.
@@
  • A GNN-based tracker to perform multi-object tracking with 3D bounding boxes in BEV coordinates with impressive long-term association capability.
  • We achieve the state-of-the-art on both the AICity’24 dataset and the WildTrack dataset.
@@ 2 Related Works
Figure 2: Comparison among 3 types of MTMC tracking methods.
@@
3 BEV-SUSHI
Figure 3: The overall framework of BEV-SUSHI. First, multi-view images at frame $t$ are passed through the image backbone to obtain multi-view image features. A spatial encoder is then introduced to aggregate multi-view image features to BEV features $B_{t}$, followed by a temporal encoder to aggregate BEV features within a temporal window. A DETR-based decoder is utilized to obtain object detection results, which are in the format of 3D bounding boxes. To get reliable ReID features for the detected objects, a ReID feature extraction module is proposed, including a 2D ReID feature extractor and a 2D-3D detection association algorithm. Finally, SUSHI-3D is designed to achieve multi-object tracking in BEV to obtain the final MTMC tracking results. (SUSHI Block graphics are from [5].)
@@ 3.1 Coordinate Systems and Projection
We first define the 3D world coordinates $\mathcal{W}$ and 3D camera coordinates $\left\{\mathcal{C}^{i}\right\}_{i=1}^{V}$ for each specific scene. Besides, we also define the 2D image pixel coordinates $\left\{\mathcal{U}^{i}\right\}_{i=1}^{V}$ for each camera image. Here, $V$ is the number of camera views in the scene. The goal is to project multi-view camera information to the 3D world and do detection and tracking in $\mathcal{W}$ without further cross-view spatial association.
As shown in Fig. 1, the origin of $\mathcal{W}$ is defined around the center of the scene, lying on the ground plane. Axes $x$ and $y$ are parallel to the ground plane, and $z$ is vertically pointing up. The origin of $\mathcal{C}^{i}$ is defined at the camera center. Axis $x$ points right, $y$ points down, and $z$ points forward. A 3D point $\mathbf{x}=[x,y,z]^{\top}$ in the world coordinates $\mathcal{W}$ can then be projected to the pixel coordinates $\mathcal{U}^{i}$ in the $i$-th camera view as $\mathbf{u}=[u,v]^{\top}$. This projection can be done by
@@ 3.2 Multi-View 3D Object Detection
Towards this end, multi-view images $\left\{I^{i}_{t}\right\}_{i=1}^{V}$ at frame $t$ are input into the image backbone to obtain multi-view image features $\left\{F^{i}_{t}\right\}_{i=1}^{V}$. The spatial encoder then aggregates multi-view features $\left\{F^{i}_{t}\right\}_{i=1}^{C}$ to BEV features $B_{t}$. We adopt the spatial cross-attention (SCA) [18] for this aggregation,
@@ 3.3 Multi-View ReID Feature Extraction
Appearance features are crucial for tracking tasks, especially for cross-view tracking, since object appearances from different cameras might differ due to varied illuminations and viewing angles. The naïve way to obtain the appearance feature for each 3D detection from BEVFormer is to project the 3D bounding box to each camera image and extract ReID features by the projected box. However, unlike the 2D bounding boxes directly from 2D object detectors, the projected 3D bounding boxes are usually much larger than the actual objects in the image (see Fig. 4). This will include noisy background or other nearby objects in the target object so that affect the quality of the ReID features. Therefore, we propose a 2D-3D detection association algorithm to effectively find the best 2D bounding box for each 3D detection. For the 2D detector, we use DINO [40] with FAN-small backbone [44]. The detector is pre-trained on a subset (800k images) of the OpenImages dataset [17], pseudo-labeled by a DINO detector trained on 80 COCO classes. Then, the detector is fine-tuned on a proprietary dataset with more than 1.5 million images and more than 27 million objects for the person class.
Next, we extract ReID features for each 2D bounding box from the 2D detector. Our person-centric ReID model is based on the SOLIDER [7], a self-supervised learning framework with a Swin-Tiny Transformer [19] backbone. The model uses image crops of size 256$\times$128 as input and outputs a feature of dimension 256 enhancing memory efficiency and throughput. We pre-train the model on a proprietary dataset of 3 million unlabeled image crops. The supervised fine-tuning is performed on a collection of both real and synthetic datasets which includes characters from Market-1501 [43], AICity’24 [31] and additional propitiatory datasets totaling the object count to 3,826 identities and image count to 82k object crops. We analyze the ReID feature quality in Appendix C.
@@
We evaluate the 2D-3D detection association algorithm to ensure we have high-quality ReID features assigned to each 3D detection. This evaluation is based on the association results between ground truth 2D and 3D annotations in the AICity’24 training set. We introduce the evaluation metric, detection association accuracy, defined as the number of corrected matches divided by the number of matched ground truth 2D bounding boxes. The evaluation results are shown in Tab. 1. We can find that even in crowded hospital scenes, our algorithm can achieve above 90% accuracy, illustrating its robustness. More visualizations based on predictions are shown in Sec. B.2.
Scene Warehouse Retail store Hospital Overall
Accuracy 99.4% 95.2% 91.2% 96.9%
@@
The early multi-view aggregation of our BEV-SUSHI framework allows us to perform multi-object tracking (MOT) directly in 3D world coordinates $\mathcal{W}$. With 3D detections from BEVFormer and features from ReID model, graphs provide a natural framework to address the long-term association challenges in MTMC tracking within a tracking-by-detection setting. Leveraging its unique long-term modeling capabilities, we build upon SUSHI [5], originally designed for 2D tracking, to introduce the first hierarchical NN-based tracking solution in 3D space.
@@
The generalizable and versatile graph structure as well as node and edge feature design in [5] facilitates its application to 3D space. For BEV-SUSHI, we replace the 2D geometry, appearance, and motion feature encodings with MTMC 3D counterparts.
@@
In contrast to 2D bounding boxes, geometric distances in 3D are not impacted by projection distortions and camera distance scaling. Therefore, we omit distances between object dimensions for our MTMC target datasets, which primarily focus on person tracking.
Our BEV-SUSHI framework extracts appearance information in the form of multi-view ReID features obtained via 3D-2D projection and detection association (described in Sec. 3.3). To provide our graph model with view-consistent and stable appearance distances, we compute cosine distances between ReID features averaged across all cameras in which the associated 3D detection is observable.
@@
AICity’24 [31] evaluates identity consistency across sequences with up to 24k frames including occlusions ranging up to 2k frames. To tackle these challenges, we train a GNN with 9 layers to associate objects for up to 3,840 frames (empirical memory limit for a 32 GB GPU) at once and process the entire sequence in a near-online sliding window fashion with stride $s=1920$. The original SUSHI [5] inference computes overlapping graph outputs and associates tracks via handcrafted, and therefore error-prone, matching heuristics.
In this work, we propose a novel near-online inference of a trained GNN hierarchy that removes potential biases in tracking heuristics and unlocks occlusion handling beyond the number of frames per graph. As depicted in Fig. 5, we omit the graph overlap and process only the incoming frames with the model hierarchy. An additional global block associates any previous track with the predictions on the incoming frames. In practice, we repurpose the last hierarchy level of our model and replace the inference of the graph overlap from previous layers with past tracks. Empirically, we observed that the layer trained for inference of 3,840 frames generalizes to predict more long-term associations without any retraining.
Figure 5: The tracking inference of BEV-SUSHI. We process long sequences in a near-online fashion with stride $s$. But in contrast to [5], we omit overlaps between windows and process only incoming frames with our graph neural network (GNN) model. To associate incoming with past tracks, BEV-SUSHI extends the input time span of the final hierarchy layer to function as a global merging block. The global block requires no additional training.
@@ 4.1 Datasets and Implementations
is an MTMC tracking benchmark consisting of 6 different synthetic environments, e.g., warehouses, retail stores, and hospitals, developed using the NVIDIA Omniverse Platform [22]. The dataset includes 90 scenes, 40 for training, 20 for validation, and 30 for testing, with a total of 953 cameras, 2,491 people, and over 100 million bounding boxes. Besides, we also include 30 additional scenes with similar camera settings for BEVFormer training. In this dataset, persons are annotated as 3D bounding boxes with 3D locations, dimensions, and orientations.
@@
is a real-world MTMC tracking benchmark containing a single sequence of a scene where 7 synchronized cameras cover over a 36$\times$12 m space at a 1920$\times$1080 resolution. There are 400 fully annotated frames per camera at 2 FPS with a total of 313 identities and 42,721 2D bounding boxes. As for the 3D annotations, there is no 3D bounding box available. A grid is defined on the ground plane and all persons are annotated by the corresponding grid indices, which can be mapped to 3D locations on the ground plane in meters.
Following the experiment setting in [11, 8, 29], we use the first 360 annotated frames as the training data and evaluate the rest 40 frames. Since there is no 3D bounding box annotation in the WildTrack dataset, we remove the bounding box dimension and orientation regression losses during BEVFormer training. We use a default 3D bounding box dimension (width, length, height) of $[0.6,0.6,1.7]$ and zero rotation for later 2D-3D detection association.

### v3

# MCBLT: Multi-Camera Multi-Object 3D Tracking in Long Videos

@@ Abstract
Object perception from multi-view cameras is crucial for intelligent systems, particularly in indoor environments, e.g., warehouses, retail stores, and hospitals. Most traditional multi-target multi-camera (MTMC) detection and tracking methods rely on 2D object detection, single-view multi-object tracking (MOT), and cross-view re-identification (ReID) techniques, without properly handling important 3D information by multi-view image aggregation. In this paper, we propose a 3D object detection and tracking framework, named MCBLT, which first aggregates multi-view images with necessary camera calibration parameters to obtain 3D object detections in bird’s-eye view (BEV). Then, we introduce hierarchical graph neural networks (GNNs) to track these 3D detections in BEV for MTMC tracking results. Unlike existing methods, MCBLT has impressive generalizability across different scenes and diverse camera settings, with exceptional capability for long-term association handling. As a result, our proposed MCBLT establishes a new state-of-the-art on the AICity’24 dataset with $81.22$ HOTA, and on the WildTrack dataset with $95.6$ IDF1.
@@ 1 Introduction
Detecting and tracking objects across multiple cameras is a crucial problem for 3D environment understanding, particularly in retail or warehouse settings, for various applications, including inventory management, security surveillance, or customer behavior analysis. In these use cases, a typical multi-camera system involves numerous cameras with diverse viewing angles and fields of view (FoV). While some cameras usually have overlapping FoVs, others may not share any common visual space. However, this task presents significant challenges such as occlusions, varying lighting conditions, and the need to maintain consistent object identification across different camera views. Moreover, the integration of 3D spatial information from multiple 2D views requires sophisticated algorithms to handle camera calibration errors and perspective distortions. Addressing these issues is crucial for developing robust multi-camera detection and tracking systems that can reliably operate in complex environments.
Figure 1: Comparison among three types of MTMC tracking methods. (a) conducts 2D detection separately and associates objects among different views by appearance-based ReID; (b) considers geometric constraints as well besides appearance for cross-view association; (c) achieves multi-view association in early stage by feature-level aggregation.
@@
Existing MTMC tracking methods can be divided into three categories: (i) late multi-view aggregation, (ii) late multi-view aggregation with geometric projection, and (iii) early multi-view aggregation, as shown in Fig. 1. Late multi-view aggregation pipelines detect objects in each camera view as 2D bounding boxes, with some methods tracking these 2D detections separately before associating them across cameras using re-identification (ReID) embeddings and spatial constraints. Late multi-view aggregation pipelines with geometric projection pipelines [35, 36, 8] further perform additional projections using camera calibration matrices to achieve spatial association for each camera view into a global coordinate system. Recently,  [30] demonstrated the possibility of early aggregation before any perception steps in each camera view. This method first fuses multi-view image information into a unified 3D space and directly conducts both detection and tracking in this 3D space, which can significantly improve detection quality and avoid association errors due to unreliable spatial alignment among different camera views. However, its detection network is designed for a fixed multi-camera scene, so it is not flexible for different environments, different numbers of cameras, or different camera placements. Therefore, it does not have good scalability for large scenes with a large number of cameras. Furthermore, the tracking of [30] merely relies on a heuristic Kalman filter, which is prone to drifting and struggles with long-term occlusion handling.
@@
In this work, we propose MCBLT, a robust and generalizable method for Multi-Camera Bird’s-eye-view Long-term Tracking in various complex multi-camera indoor/outdoor public environments. Our proposed approach leverages a unified BEV representation and spatio-temporal transformers to directly infer 3D object bounding boxes. Then, the well-localized 3D detections are backprojected to 2D camera images and matched with 2D detections for optimal 2D appearance feature extraction. This allows us to benefit from both highly accurate 2D and 3D detections for ReID and long-term tracking. For the temporal association of detections, we introduce the first 3D tracking-by-detection approach utilizing hierarchical graph neural networks (GNNs) [5]. The lightweight hierarchical structure allows us to learn a GNN tracking model that directly predicts associations for up to thousands of frames. Compared to prior 2D GNN tracking solutions [5], we scale the ability to bridge occlusion gaps by an order of magnitude. Our model consists of a hierarchy of GNNs that associate 3D detections via geometrical and multi-view ReID embedding features. To track objects across very long sequences, we introduce a novel global tracking layer that replaces handcrafted heuristics required to run the GNN in a sliding window fashion [5] with a model layer that operates globally over the video sequence and matches incoming objects with past tracks. The global layer requires no additional training and significantly boosts occlusion handling and overall tracking performance.
@@
We evaluate MCBLT on a large-scale synthetic indoor MTMC dataset AICity’24 [32] and a real-world outdoor MTMC dataset WildTrack [6]. MCBLT achieves SOTA on both datasets, with $81.22$ HOTA on AICity’24, and $95.6$ IDF1 on WildTrack.
@@
  • The first MTMC method efficiently performs early multi-view image aggregation for 3D perception.
@@
  • The first hierarchical GNN-based 3D multi-object tracking method and a novel global tracking block that unlocks long-term associations across thousands of frames.
  • We achieve state-of-the-art performance on both the AICity’24 and WildTrack datasets.
@@ 2 Related Works
3 MCBLT Framework
Figure 2: The overall framework of MCBLT. First, multi-view images at frame $t$ are passed through the image backbone to obtain multi-view image features. A spatial encoder is then introduced to aggregate multi-view image features to BEV features $B_{t}$, followed by a temporal encoder to aggregate BEV features within a temporal window. A DETR-based decoder is utilized to obtain object detection results, which are in the format of 3D bounding boxes. To get reliable ReID features for the detected objects, a ReID feature extraction module is proposed, including a 2D ReID feature extractor and a 2D-3D detection association algorithm. Finally, SUSHI-3D is designed to achieve multi-object tracking in BEV to obtain the final MTMC tracking results. (SUSHI Block graphics are from [5].)
@@ 3.1 Coordinate Systems and Projection
We first define the 3D world coordinates $\mathcal{W}$ and 3D camera coordinates $\left\{\mathcal{C}^{i}\right\}_{i=1}^{V}$ for each specific scene. Besides, we also define the 2D image pixel coordinates $\left\{\mathcal{U}^{i}\right\}_{i=1}^{V}$ for each camera image. Here, $V$ is the number of camera views in the scene. The goal is to project multi-view camera information to the 3D world and perform detection and tracking in $\mathcal{W}$ without further cross-view spatial association.
The origin of $\mathcal{W}$ is defined around the center of the scene, lying on the ground plane. Axes $x$ and $y$ are parallel to the ground plane, and $z$ is vertically pointing up. The origin of $\mathcal{C}^{i}$ is defined at the camera center. Axis $x$ points right, $y$ points down, and $z$ points forward. A 3D point $\mathbf{x}=[x,y,z]^{\top}$ in the world coordinates $\mathcal{W}$ can then be projected to the pixel coordinates $\mathcal{U}^{i}$ in the $i$-th camera view as $\mathbf{u}=[u,v]^{\top}$. This projection can be done by
@@ 3.2 Multi-View 3D Object Detection
Towards this end, multi-view images $\left\{I^{i}_{t}\right\}_{i=1}^{V}$ at frame $t$ are input into the image backbone to obtain multi-view image features $\left\{F^{i}_{t}\right\}_{i=1}^{V}$. The spatial encoder then aggregates $\left\{F^{i}_{t}\right\}_{i=1}^{V}$ to BEV features $B_{t}$. We adopt the spatial cross-attention (SCA) [18] for this aggregation,
@@ 3.3 Multi-View ReID Feature Extraction
Appearance features are crucial for tracking tasks, especially for cross-view tracking, since object appearances from different cameras might differ due to varied illuminations and viewing angles. The naïve way to obtain the appearance feature for each 3D detection from BEVFormer is to project the 3D bounding box to each camera image and extract ReID features by the projected box.
However, the projected 3D bounding boxes are usually much larger than the actual objects in the image (see Fig. 3). This will include noisy background or other nearby objects in the target object so that affect the quality of the ReID features. Therefore, we propose a 2D-3D detection association algorithm to effectively find the best 2D bounding box for each 3D detection.
@@
The early multi-view aggregation of our MCBLT framework allows us to perform multi-object tracking (MOT) directly in 3D world coordinates $\mathcal{W}$. Graphs provide a natural framework to address the long-term association challenges in MTMC tracking within a tracking-by-detection setting. We build on SUSHI [5], originally designed for 2D tracking, to introduce the first hierarchical GNN-based tracking solution in 3D space. To this end, we first extend its graph formulation and features to 3D. Then we tackle the weaknesses of SUSHI’s long-term tracking, i.e., heuristic matching of overlapping windows and occlusion handling being limited to the window size by introducing a novel global block to the hierarchy. Overall, we demonstrate long-term tracking on a previously unseen temporal scale (SUSHI: 512 frames vs. MCBLT full sequence).
@@
The generalizable and versatile graph structure as well as node and edge feature design in [5] facilitates its application to 3D space. For MCBLT, we replace the 2D geometry, appearance, and motion feature encodings with MTMC 3D counterparts.
@@
In contrast to 2D bounding boxes, geometric distances in 3D are not impacted by projection distortions and camera distance scaling.
MCBLT extracts appearance information in the form of multi-view ReID features obtained via 3D-2D projection and detection association (described in Sec. 3.3). To provide our graph model with view-consistent and stable appearance cues, we compute cosine distances of ReID features averaged across all cameras in which the object is observable.
@@
Figure 4: GNN hierarchy of MCBLT for tracking. We process long sequences in a near-online fashion with stride $s$. But in contrast to [5], we omit overlaps between windows and heuristics. To associate incoming with past tracks, MCBLT uses a global merging block. The global block requires no additional training.
@@
Due to their larger camera coverage, multi-camera setups can observe objects across longer time spans and, therefore, pose uniquely challenging long-term tracking problems. For example, AICity’24 [32] consists of sequences with up to 24k frames including object occlusions for up to 2k frames, a magnitude longer compared to popular single-view tracking benchmarks [21] The original SUSHI [5] inference computes overlapping graph outputs and associates tracks via handcrafted and, therefore, error-prone matching heuristics.
In this work, we propose a novel near-online inference and global tracking block that associates previous tracks with predictions on the incoming frames. Figure Fig. 4 depicts our new inference without graph overlaps. Our GNN hierarchy first processes the non-overlapping incoming frames with regular SUSHI blocks and then associates objects across the full sequence with the final global tracking block. We process the entire sequence in a sliding window fashion.
The global block shares weights with the previous hierarchy level and thus requires no additional training. Our MCBLT removes potential biases in previous heuristics in GNN-based tracking and unlocks occlusion handling beyond the number of frames per graph.
@@ 4.1 Datasets
is an MTMC tracking benchmark consisting of 6 different synthetic environments, e.g., warehouses, retail stores, and hospitals, developed using the NVIDIA Omniverse Platform [23]. The dataset includes 90 scenes, 40 for training, 20 for validation, and 30 for testing, with a total of 953 cameras, 2,491 people, and over 100 million bounding boxes. Besides, we also include 30 additional scenes with similar camera settings for BEVFormer training. In this dataset, persons are annotated as 3D bounding boxes with 3D locations, dimensions, and orientations. Long-term tracking performance is crucial for AICity’24 [32] as it evaluates identity consistency across sequences with up to 24k frames including occlusions ranging up to 2k frames.
@@
is a real-world MTMC tracking benchmark containing a single sequence of a scene where 7 synchronized cameras cover over a 36$\times$12 m space at a 1920$\times$1080 resolution. There are 400 fully annotated frames per camera at 2 FPS with a total of 313 identities and 42,721 2D bounding boxes. As for the 3D annotations, there is no 3D bounding box available. A grid is defined on the ground plane and all persons are annotated by the corresponding grid indices, which can be mapped to 3D locations on the ground plane in meters. Following the experiment setting in [11, 8, 30], we use the first 360 annotated frames as the training data and evaluate the rest 40 frames.
4.2 Implementation Details